40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 21 15:13:44 2017, gautam, Update, IMC, IMC input beam mode matching IMG_7220.JPGIMC_ModeMatch.pdfsingleLensSensitivity.pdfsensitivity.pdfIMCmodeMatch.m
    Reply  Tue Mar 21 21:59:48 2017, gautam, Update, IMC, IMC input beam mode matching IMG_3654.JPGIMG_3655.JPG
       Reply  Wed Mar 22 00:33:00 2017, gautam, Update, IMC, IMC length offset nulling offsetInvestigation.pdfoffset_summing_amp.pdf
          Reply  Wed Mar 22 16:58:25 2017, gautam, Update, IMC, WFS sensing matrix measurements 
             Reply  Thu Mar 23 01:44:53 2017, gautam, Update, IMC, WFS sensing matrix measurements IMC_WFS_segment_TF.pdfIMC_WFS_channels_TF.pdfTFsummary.pdfIMC_WFS_170322.xlsx.zip
                Reply  Thu Mar 23 08:43:11 2017, rana, Update, IMC, WFS sensing matrix measurements 
                   Reply  Thu Mar 23 23:38:58 2017, gautam, Update, IMC, MC SUS damping gains stepped down 
                      Reply  Fri Mar 24 11:26:57 2017, gautam, Update, IMC, MC SUS damping gains restored 
                         Reply  Fri Mar 24 19:04:18 2017, gautam, Update, IMC, Seismic feedforward and WFS MCL_comparison.pdfseis_comparison.pdf
Message ID: 12899     Entry time: Wed Mar 22 00:33:00 2017     In reply to: 12898     Reply to this: 12900
Author: gautam 
Type: Update 
Category: IMC 
Subject: IMC length offset nulling 

[valera, gautam]

Motivation: see this elog

I was fiddling around for a few days trying to implement the method outlined in this paper to null this offset - I will post a separate elog about my efforts but Valera pointed out that we could try injecting an AF modulation at the IN2 input of the MC Servo Board. Last night, we hooked up an SR function generator (f = 312Hz, A = 0.01Vpp, IN2 gain = -5dB) to the unused BNC IN2 input of the MC Servo board. To avoid any additional offsets from the AO path during this measurement, I disconnected the LEMO cable (it is labelled).

We looked at the spectrum of the MC transmission around 312Hz and also 2*f = 624Hz. As a result of this modulation, we expect in the transmitted power, dP/P, a 2f term with amplitude ~(X_mod/X_0)^2 and a term at f with amplitude ~(X_offset * X_mod / X_0^2) - I may have missed out some numerical factors of order 1. So the latter should vanish if the offset at the error point is truly zero and the lock-point is the center of the resonance. Last night, we found that an offset in the range of -0.25 V to -0.19 V nulled this peak in the DTT spectrum. Today, the number was -0.05V. So the true offset seems to vary from lock to lock. Here are spectra around f=312Hz for a few different values of the offset slider (the center of the resonance seems to be -0.05V on the MEDM slider at this time).

Do these numbers make sense? Some time ago, I had pulled out the MC Servo board to find out what exactly is going on at this offset summing point. The MEDM slider goes from -10V to 10V, and by measuring the voltage at TP5 (see schematic below), I found that there is a 1/40 scaling factor between what is actually applied and the number on the MEDM slider (so for example, the numbers in the legend in the above plot have to be divided by 40). I've modified the MC Servo Board MEDM screen to reflect this. When I had pulled the board out, I noticed that in addition to the offset voltage applied via the backplane connector, there was also a potentiometer (R50 in the schematic below). I had nulled the voltage at TP5 using this potentiometer, but I guess drifts of ~5mV are possible. 

Discussion on calibration of offset slider in Hz/V:

I've yet to do a rigorous calibration of this slider into Hz, but looking at the spectrum of the transmitted intensity at 2f, we estimated the coefficient (X_mod/X_0) ~ 3e-3 for an offset of 0.2V. dP/P ~1 when the applied modulation equals the linewidth of the cavity, which is 3.6kHz. So 0.2V of offset slider corresponds to ~ 10Hz frequency offset. In other words, I estimate the slider calibration to be 50Hz/V. So with the full range of +/- 10V, we should be able to scan ~1kHz of frequency offset. What does this imply about the variation of the offset slider value that removes the peak at 1f between locks? As mentioned above, this variation is ~0.2V over a day - with the calibration mentioned above, this corresponds to a change in cavity length of ~10um, which seems reasonable to me...


So how did all of this tie in with WFS SUM offsets? We did the following:

  • After nulling the length offset using the procedure detailed above, we noticed non-zero offsets on both WFS1 and WFS2 "I" SUM outputs
  • So we set the dark offsets and RF offsets for the WFS, with no light incident on the WFS (PSL shutter closed). 
  • Re-locking the IMC and closing the WFS loops, we noticed that WFS2 SUM offset was still hovering around 0, but WFS1 SUM offset was ~ -2000cts.
  • Looking at some trends on dataviewer, this offset seems to drift around over a few days timescale by a few thousand counts - for example, the WFS1 offset today was +2000cts. Moreover, the WFS1 offset seems to drift around by ~factor of 3 times as much as WFS2 offset in the 24 hour period I looked up (plot to follow)...
  • Misaligned MC2 and looked at the sum offset with just the single bounce beam off MC1 onto the WFS

I neglected to screenshot the StripTool from the times we were doing these trials but I have the times, I will pull up some dataviewer plots and upload them here tomorrow...

Attachment 1: offsetInvestigation.pdf  23 kB  | Hide | Hide all
offsetInvestigation.pdf
Attachment 2: offset_summing_amp.pdf  75 kB  | Hide | Hide all
offset_summing_amp.pdf
ELOG V3.1.3-